OpenStudy (anonymous):

|x – 6| = 8 PLEASE HELP!

OpenStudy (anonymous):

well x = -2 here :D

OpenStudy (anonymous):

cuase there are | | there it means the process'es result will always be positive (sorry for my bad english)

OpenStudy (anonymous):

There will ALWAYS be two answers to any absolute value equation of this type

OpenStudy (anonymous):

x=14 , x=-2

OpenStudy (stormfire1):

x = -2 or x= 14 :)

OpenStudy (anonymous):

should be -2<x<2 i thnk

OpenStudy (anonymous):

i just woke up...someone needs to double check that

OpenStudy (anonymous):

|x-6|=8 by definition of absolute value becomes x-8=8 or -(x-6)=8

OpenStudy (anonymous):

well x can be 14 too :D err amistre can you help me on my question? :D which is at left :) (shared over 5 or 6 hard calculus questions today :) really they are not questions that can be wasted :) )

OpenStudy (anonymous):

x-6 = 8 or - (x-6) = 8

OpenStudy (anonymous):

x = 14 or -2

OpenStudy (anonymous):

korcan ur question is a beat

OpenStudy (anonymous):

the possible answers are: 14 or 2, 14, 14 or –2, or 14 or –14

OpenStudy (anonymous):

Many people just translate it to\[x-6=\pm8\]\[x=6\pm8=-2 \ or \ 14\]

OpenStudy (anonymous):

the answer is definitely 14 or -2

OpenStudy (anonymous):

k, thank you!